Solve the following equations and check your results. 
3m = 5m – 8/5 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

3m=5m–8/5
⇒5m–3m=8/5
⇒2m=8/5
⇒2m×5=8
⇒10m=8
⇒m=8/10
⇒m=4/5
Putting the value of m in RHS and LHS we get,
⇒3×(4/5)=(5×4/5)–8/5
⇒12/5=4–(8/5)
⇒12/5=(20-8)/5
⇒12/5 = 12/5
⇒LHS=RHS

Thanks